OpenStudy (anonymous):

*** WILL MEDAL *** Which sequences are geometric? Check all that apply. –2.7, –9, –30, –100, ... –1, 2.5, –6.25, 15.625, ... 9.1, 9.2, 9.3, 9.4, ... 8, 0.8, 0.08, 0.008, ... 4, –4, –12, –20, ...

OpenStudy (xmissalycatx):

In mathematics, a geometric progression, also known as a geometric sequence, is a sequence of numbers where each term after the first is found by multiplying the previous one by a fixed, non-zero number called the common ratio.

OpenStudy (anonymous):

I dont understand the non zero part :(

OpenStudy (anonymous):

Is the 4th one wrong? @xMissAlyCatx

OpenStudy (xmissalycatx):

I do believe so! And I'm sure the 2nd one is too.

OpenStudy (anonymous):

and so is the first one, right? @xMissAlyCatx

OpenStudy (xmissalycatx):

I couldn't find a pattern in it so..

OpenStudy (trojanpoem):

To find if a sequence is either Arithmetic or geometric: Geometric: (A long the whole sequence) \[\frac{ a_{n+1} }{ a_{n} } = const\] Arithmetic: (A long the whole sequence) \[a_{n+1} - a_{n} = const \] So walk through each sequence, like so: 1) \[\frac{ a_{2} }{ a_{1} } = \frac{ -9 }{ -2.7 } = \frac{ 10 }{ 3 } , \frac{ a_{3} }{ a_{2} } = \frac{ -30 }{ -9 } = \frac{ 10 }{ 3}\] (The first sequence is geometric) Can you continue ?

OpenStudy (xmissalycatx):

Thank you @trojanpoem